Calculate the formula mass for the following:

a) Cd3(PO4)2

b) aluminum sulfate
Answer #1

Answer:

a) Cd3(PO4)2

Mass of Cd=112.411 g/mol, P=30.973 g/mol, O= 15.99 g/mol

Three cadmium atoms=3*112.411 g/mol

Two Phosphorus atoms=2*30.973 g/mol

Eight oxygen atoms=8*15.99 g/mol

The formula mass of Cd3(PO4)2=527.175 g /mol.

b) Aluminum sulphate Al2(SO4)3

Mass of Al=27.0 g/mol, S=32.07 g/mol, O=15.99 g/mol

Two atoms of Aluminium=2*27 g/mol,

Three atoms of Sulphur=3*32.07 g/mol

Twelve atoms of oxygen=12*15.99 g/mol

Total mass of Aluminium sulphate is 342.2 g/mol.